OpenStudy (anonymous):

Solve the following inequality. x^2(6+x)(x-6)/(x+2)(x-2) ≥0

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

is that \[\frac{x^2(6+x)(x-6)}{(x+2)(x-2)} \ge 0\]

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Hmm...let me think about how to do this correctly...

OpenStudy (anonymous):

ok =)

OpenStudy (whpalmer4):

Well, first thing I'll observe is that the function is undefined at \(x = \pm 2\) because the denominator becomes 0 at those two values.

OpenStudy (whpalmer4):

Similarly, it's going to be 0 at \(x =0,\,x=\pm 6\) because the numerator is 0 at those values of \(x\). Agreed?

OpenStudy (whpalmer4):

From x = 2 and greater, the denominator is positive.

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

Now, we need to figure out what range for x=2 and greater will also have a positive numerator, since to get the result >= 0, we want the numerator to be positive when the denominator is positive. If we just look at the numerator:\[x^2(x+6)(x-6)\]That's positive if \(x\ge 6\) At \(x=0\), the numerator is 0, so that's also in our solution From \(0 < x < 6\) the numerator is negative, because of the \(x-6\) product term Agreed with all of that?

OpenStudy (anonymous):

yes =)

OpenStudy (whpalmer4):

Ah, but I was wrong, and you shouldn't have agreed so quickly :-) From \( 0 \le x \le 2\) the numerator and the denominator are both negative, so that gives us a positive quotient! for example, at x = 1 \[\frac{(1)^2(1+6)(1-6)}{(1+2)(1-2)} = \frac{1*7*-5 } {3*-1 } = \frac{35}{3}\] So I believe our answer is (for the right side only) \[0 \le x \lt 2 \,|| \,x \ge 6\]Where \(||\) means or

OpenStudy (whpalmer4):

And I think this is symmetrical about the y-axis, so really our answer is \[x \le -6 \,||\,-2< x < 2\,||\,x \ge 6\]

OpenStudy (whpalmer4):

Now for the moment of truth: I plot the function and see what it looks like :-) Hey, we were right!

OpenStudy (anonymous):

it says to type the answer in interval notation?

OpenStudy (whpalmer4):

Okay, in interval notation I think that would be \[[-\infty,-6],(-2,2),[6,\infty] \]but I don't spend any time using interval notation, so I make no promises...

OpenStudy (anonymous):

looks about right lol

OpenStudy (whpalmer4):

I'm more concerned with finding the right answer, and less with jumping through what I regard as artificial hoops (you solved the problem correctly, but got it wrong because you didn't type it the way I wanted it typed)

OpenStudy (anonymous):

ya its dumb.. i agree. can I get your help with just one more, and then i will leave you alone? lol

OpenStudy (whpalmer4):

Unfortunately, I have to go run some errands, but I'll be back later on in the evening, I think.

OpenStudy (anonymous):

oh ok =) thank you so much for your help!!

OpenStudy (whpalmer4):

You're welcome!